OpenStudy (anonymous):

parametrize the curve y=ln(2x+1) using its slope m as the parameter ..

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

slope of curve is its derivative, y' so: y' =2/(2x+1) ------>2x+1 = 2/y' puting this in the equation: y=Ln(2/y') you have your result

OpenStudy (anonymous):

answer: x=(2-m)/2m y=ln[2/m)

OpenStudy (anonymous):

just name y' = m

OpenStudy (anonymous):

just keep in mind that m(x)

OpenStudy (anonymous):

hm ok thanks myko